I have a Final exam tommorow please help. Question: This width of a rectangle is 2cm less than it's length. If the area of the rectangle is 168 cm2 (squared), what is it's width?
i think it's base x height so like x (x-2) = 168 ... x2(sqd) - 2x=168 but then what?!

Answer by checkley71(8403)

x(x-2)=168 or x~2-2x-168=0 or (x-14)(x+12)=0 or x=14cm for the length so the width is 14-2 or width is=12cm.